OpenStudy (anonymous):

plss help x=2 and y=3 ...the question is 3xy divide y^2 if i not reply im afking but plss help

OpenStudy (anonymous):

\[\frac{3xy}{y^2}=\frac{3x}{y}\] if x = 2, y = 3 you get \[\frac{3\times 2}{3}=2\]